VISTA IOT MINUTES, 2010-01-26, 14-16h Garching / 13-15h UK / 10-12h Chile (prepared by Magda Arnaboldi, imput from VIV on PPRS section) * * * Present: Garching: Marina Rejkuba (MRE), Magda Arnaboldi (MAR), Wolfgang Hummel (WHU), Michael Hilker (MHI), Lander de Bilbao (LDB), Gaitee Hussain (GHU). Santiago (via phone): Thomas Szeifert (TSZ) Paranal: Valentin D. Ivanov (VIV), Elena Pompei (EPO) UK (via phone): Jim Emerson (JEM), Jim Lewis (JLE), Mike Irwin (MIR), Monika Petr Gotzens (MPE) * * * Agenda: 1. VISTA status and planned activities (TSZ) 2. Phase II progress (MRE) 3. VISTA PPRS Summary (VIV) 4. Science Verification Data Processing Report (MAR, MRE, MPE, VIV) Will we be ready with a paper draft for the next Messenger? 5. Action Items review (see bellow a summary from the 2009-12-01 IOT) 6. Others 7. Date of the next meeting: end of Feb, 2010, possible mini-IOT before the start of the next "period" * * * 1. VISTA status, schedule and planned activities TSZ: current negotiations with consulting engineering company about M2 and enclosures. Current plan includes possible shut-down of the telescope for a couple of days. On 03/02//2010 there is a a visitor mode, and after that there is going to be the installation of the new VIRCAM templates, and the preparation for phase 2. TSZ will ask the TIO to look into the current master flats, as their current quality may be affected by saturated frames. TSZ asks WHU to send a new library of flats and master calibrations for installation in the pipeline workstation. MRE asks whether there are plans to make the night reports and info on OBs execution an quality flags to PIs/CASU. TSZ reports work in progress by Cecilia. It requires the restructuring of the database. It will happen while JEM is on Paranal. It will require 2-3 weeks more work. Currently there are problems in formatting the nightlog. WHU asks whether the current VISTA nightlog are ingested in the Paranal database. TSZ replies that they are not, only the headers are ingested but there are no inputs on the single OBs. VIV comments that only the comments which are relevant for the telescope operations are ingested. TSZ data on the time execution of OBs are ingested in the database but a tool is required to extract this information and make it available. This is what Cecilia is doing, and requires 2-3 weeks. VIV reports that they had a Visitor mode run at VISTA, and JEM will observe starting from Feb 3, 2010. JEM asks whether the same OB can be executed more time in visitor mode, and TSZ stated yes. * * * 2. Phase II progress (MRE) MRE: the deadline for the PS OB submission is Friday, February 22, 2010. Only 50% of the Obs were submitted. VMC will submit OBS in a month as the LMC is observable only at the end of August. MHI: reports one full submission only. Both VHS and VIKING are late and they should submit before Feb 15, 2010. TSZ reports of a difficult execution of the VHS obs for dry runs, as these OBS were submitted late, and the target field was setting. Obs were executed by artificially boosting their priorities . TSZ reports that Thomas Bierwirth will be on Paranal on Feb. 8, 2010 for the installation of the new OT, and the IP change. MRE informs that the current Phase 2 submission covers the time interval from Feb 15 to end September. The next Phase 2 submission for PS OBs will phase in with the VLT normal programs. So the next deadline for Phase 2 submission will be in the first half of August, 2010. The Phase 2 is normally open for 6 weeks, from end of June to beginning of August. USD will need the final version of SADT by end of June. JEM & MRE & MHI: expected delivery of final SADT by end of April, interactions ESO USD & JEM, acceptance meeting in June. MHI: regarding SADT, ESO will not provide the extension of the 2MASS catalog to z. MHI suggest further discussion of ESO/JEM for the implementation of the new parameters for the use of the UCAC3 catalog. 3. VISTA PPRS Summary (VIV) VIV reports that there are 62 open PPRS's. The most common are for the auto-guiding failures and for various software issues. By assignment: - 3 for Computer Hardware, mostly about larger disks - 12 for Electronics, dominated by the M2 hexapod, various lamps - 4 instrumentation; the most serious are the Detector No. 16 upper part being sub-sensitive, and the filter wheel initialization - 2 for Maintenance: alarms, workstation reboots - 4 for Mechanical: mostly M1 and M2 (we have many PPRS's for the M2 hexapod) - 1 for Safety: about the radio - 11 for Science Operations: typically for various confirmations, tests, Remedy, OT, and various items from the punch list - 25 for Software: mostly for AO and AG problems About 55 are PPRS related to the telescope and they are usually about auto-guiding, M1, and M2 problems. Only few of the tickets deals with the instrument/detector and they are about Detectors 14 (the infamous channel 11) and 16, a suspected inaccuracy of the filter wheel initialization; a few camera software issues (i.e. the ICS redatum always produces an error message but actually does initialize the wheel). PPRS breakdown by status: - 6 new - 41 assigned - 6 closed VIV: About detector #16: about 1/3 has low QE. VIV thinks this is a problem with the detector, not the electronics. JEM: says that Det. 16 has always been like this. JLE states that QE is OK in H and K bands, but worse in the bluer bands. TSZ reports pick-up noise in channel 11 of detector #14. JLE confirms that additional noise is visible in the raw data and in the reduced paw prints. TSZ states that one needs to understand the hardware problems that generates the pick-up noise in the electronic signal. Discussion developed whether changing the boards will solve the problems. VIV points that changing a board will affect a lot of parameters, and not one detector only, but typically in 4, and asks for more information from CASU, e.g. when the noise problem started and how frequent it is. JLE replies that it may be difficult to pin it down as it can be barely seen while doing imexam with IRAF. It looks like the bias is slightly off. Conclusions are that that Science Operations should not worry too much about it, but monitor the effect and act if it worsens. JLE explains that this is _much_ weaker than the noise that was experienced in Channel 11 on Detector 14. WHU: about Detector #16, the parameters for the upper parts of this detector turned out to be stable during the last months (i.e. on a month scale). Nonetheless, the noise itself changes from pixel to pixel very quickly. In other words, the detector is stable on large scale, but not on small pixel-to-pixel scale. 4. Science Verification Data Processing Report (MAR, MRE, MPE, VIV) MPE: CASU reduced data (stacked pawprints) for the Orion mini-survey are now in Garching on vistasv1 since Mid Dec. Access to the Orion data at VSA was provided to Orion SV team members. E. Hatziminaoglou has carried out 1st order check on quality of the images on NGC 253. T. Stanke has worked on the molecular clouds, Simon (CASU) has produced the band merged catalogs. Check on astrometry on-going. Current works on the 1st catalog, improved on the second version. Simon currently working on the variability of Tstars. VIV asked about the comparison with the photometry of 2MASS. he reported a weak gradient in J, below 0.1. MIR states whether the correct color equation was used, because the J band in the catalog is not 2MASS J band but VISA J band, so a color equation should be used. VIV asks for details on the color equation - MIR available from CASU webpages. MAR reported on current developments for the NGC 253 and work in progress for the ESO messenger paper on the VISTA SV. MPE ask for the outline to be circulated so that she can add the contribution for Orion. 5: Action Items Progress report on previous action items: * AI 2009-01: Monitor execution times & survey completion status Is being followed up by MAR, MRe. 2009-07-29: There is a first prototype 2009-10-06: no news (MRe) 2009-12-01: no news 2010-01-26: no news STATUS: ONGOING * AI 2009-04: JEm: get in contact with F. Comeron to formalize the SADT delivery. Not yet ready to do this, though JEm chatted to Fernando about it. Documentation still missing. Discussion about the magnitude ranges that SADT takes for AG/AO stars. TSz: this is a configuration issue. Thought to be ready On September 6, 2009 2009-10-06: a report will be ready on Nov 17 (MAr) 2009-12-01: the report is now ready (MAr) 2010-02-26: agreed plan for the delivery of final version of SADT by end of April, and acceptance review in June STATUS: ONGOING * AI 2009-06: VIV: by the end of June send the new simulation report and code to be taken over for future simulation. 2009-10-06: report writing in progress (VIv); to be transferred to the EST during Oct SciVer (TSz) 2009-12-01: MRe: the simulation delivered to OPO and Gaitee takes over. VIv: the completeness rate is ~85%, we will need one extra year to complete the already approved surveys. TSz: we have to look again at the overheads 2010-01-26 GHU reports on her advancement with VIV's code. More work to be done though. Further complications may arise form the request of scheduling Chilean time which is not allocated to VVV. and from open use programs in P88. STATUS: ONGOING * AI 2009-11: TSz should find out if the constraints sets could be carried with the OBs, to the OS and the fits header of the VIRCAM files. Otherwise it is not easy to propagate these data by interfaces. 2009-07-29: TSz iterating with R. Schmutzer, but DFI (T. Bierwirth) will also need to be involved. For surveys, given the small range of programs it is expected that there are not too many different OBs, and this might need to be followed in a different/manual way in the beginning - this could imply a major change for the tools that cannot be done in the last moment. STATUS: ONGOING * AI 2009-12: TSz should check how/if the fits keywords of the raw file have the information about the concatenation and other scheduling containers. 2009-07-29: Answer is NO. But is this necessary? JEm thinks yes if processing of data takes this into account. TSz: this will need to be done together with P2PP/OT programmers and VLT software support after discussion. TSz will raise it, but cannot promise a solution. 2009-10-06: the information about the containers must be included in the keywords, CASU must be included in this work because it will have an effect on the pipeline (JEm); this is also relevant for the sky offset template (MAr); not really, because the sky offset is in the same OB as the primary target tile (JEm); in the past we used for this purpose the DPR keywords (TSz); this info might be needed also for the ESO archive users (WHu); a good example for the archive is to identify the telluric standards, corresponding to a given data set, so the issue is relevant generally for all ESO instruments (MRe). 2009-12-01: MRe and Th. B. are looking into it, working with the BoB people. STATUS: ONGOING * AI 2009-14: GHu to check that the seeing distribution function in the scheduler for VIRCAM takes into account the finite (0.51") nominal resolution of the instrument. 2010-01-26: GHu states that the seeing distribution function can be implemented any moment, but it has not be done yet. STATUS: ONGOING * AI 2009-20: installation of a new server for faster/cheaper/better data reduction, to be available for semi-real time data reduction during the Science Verification run on 16-30.10.2009. TSz will talk to Christophe Dumas. 2009-12-01: the data had to be compressed to ensure that a copy is kept until the USD disks arrive in Garching and are checked 2010-01-26 most problems are related to disk space (TSZ) STATUS: CLOSED * AI 2009-21: Create poor weather conditions (particularly poor seeing) back up SciVer programs. The approved SciVer programs request seeing 0.8 arcsec or better, if the SciVer was last week, it would result in almost zero data. STATUS: CLOSED * AI 2009-22: Increase the jitter box size of the sky flats to ~1/2 of the detector size. 2010-01-26: The jitter box was increased to +/-300 arcsec. JLE confirmed that this should be fine. STATUS: CLOSED * * New Action Items from this meeting: * AI 2009-23: JEm should deliver the SADT by Dec 10 STATUS: CLOSED * AI 2009-23: MHi will contact Andine about extrapolating the 2MASS to Z-band magnitudes. 2010-02-26 MHI reported the negative answer STATUS: CLOSED AI 2010-01: TSZ the calibration database should be updated with new flats and master calibrations and how to do it regularly, as no procedure is currently in place. STATUS: NEW - WHU should give Thomas Bierwirth a usb disk with the updated calibrations before Feb7, 2010. Then on try to define a procedure for a regular update of reference frames on Paranal AI 2010-02: TSZ: how many standard stars fields should be taken? Effort in February to take up to 3 fields. Request to implement a change of jitter pattern by JLE, from 4 to 5 jitter. Try to keep calibrations plan withing 1 hrs total execution time. STATUS: NEW - JLE suggests to discuss issues with PIs, and provide feedback to TSZ, especially in the ZY band calibrations in particular. * * * 6: Others N.A. * * * 7: Date of the next meeting: 16 Feb, 2010, 14 h Garching / 10 h S-go a short "readiness meeting" on Feb 10th for USD and SCIOPs